OpenStudy (anonymous):

What is the slope of the line through the points (−7,3) and (−5,4)?

OpenStudy (anonymous):

The slope is the rise over the run - the change in y over the change in x.

OpenStudy (anonymous):

like they said, \[\frac{4-3}{ -5-(-7)}=\frac{ 1 }{ 2}\]